Problem 3:

Astro Corporation has budgeted sales and production over the next quarter as follows:

July

August

September

Sales in units

44,500

56,500

?

Production in units

45,250

56,800

61,150

The company has 4,900 units of product on hand at July 1. 10% of the next month's sales in units should be on hand at the end of each month. October sales are expected to be 76,000 units.

Required: Budgeted sales for September would be (in units)?
Answer #1

Calculation of August closing stock:

Opening stock of August = 10% X 56,500 = 5,650

Closing stock = Production in units + Opening stock - Sales in units

= 56,800 + 5,650 - 56,500

= 5,950

Closing stock of August is equal to 10% of September sales.

September sales = 5,950 / 10% = 59,500